Hi, I need help with a problem involving systems of linear equations, addition (elimination)method. The answer to the problem is ( 2,1) according to the back of my text book but everything I have tried has resulted in different answers. I would greatly appreciate your help with this one and Thanks.
Here's the problem:
3x-4y=2
4x+3y=1

Answer by stanbon(75887)

The answer to the problem you posted is definitely not (2,1)
-----------------------------
3x-4y=2
4x+3y=1
-------------
Multiply thru the 1st equation by 3; Multiply thru the 2nd eq. by 4:
----------------------------
9x - 12y = 6
16x + 12y = 4
-----------------
25x = 10
x = 2/5
-----
Substitute into 3x-4y = 6
3(2/5)-4y = 6
(6/5)-4y = 6
-4y = 6-(6/5)
-4y = (30/5)-(6/5)
-4y = (24/5)
y = -6/5
======================
Cheers,
Stan H.